OpenStudy (konradzuse):

integration of (cos(x))/(sin(x)+(sin(x))^2)

OpenStudy (konradzuse):

\[\frac{\cos(x)}{\sin(x) +(\sin(x))^2}\]

OpenStudy (anonymous):

let t=sin x

OpenStudy (anonymous):

I think: Put: \[1 + sinx = t\]

OpenStudy (anonymous):

@KonradZuse \[dt = \cos x \ dx\]

OpenStudy (anonymous):

\[\frac{cosx}{sinx(1 +sinx)}dx\]

OpenStudy (anonymous):

Put: \[1 + sinx = u\] \[sinx = (u-1)\] -------------1 \[du = cosxdx\] \[\int\limits_{}^{}\frac{du}{u(u-1)}\]

OpenStudy (konradzuse):

interesting.

OpenStudy (konradzuse):

so then do we do u^2 - u?

OpenStudy (konradzuse):

then it owuld be u^3/3 - ln|u| +c?

OpenStudy (anonymous):

Now use completion of square method..

OpenStudy (anonymous):

just notice \[\frac{1}{u(u-1)}=\frac{1}{u-1}-\frac{1}{u}\]

OpenStudy (anonymous):

Yeah mukushla is right you can use this it will become easy to solve..

OpenStudy (konradzuse):

Idk how we would do completing the square. So my method was wrong huh...

OpenStudy (anonymous):

No go by mukushula method.. He has just simplified it for you above..

OpenStudy (konradzuse):

I guess it would be ln|u-1| - ln|u| + c?

OpenStudy (anonymous):

right

OpenStudy (konradzuse):

I see what I did and why it's not right, but you couldn't just put them back together then huh...

OpenStudy (konradzuse):

ln|sin(x)| - ln|1+sin(x)| +c

OpenStudy (konradzuse):

is that correct?

OpenStudy (anonymous):

yes

OpenStudy (konradzuse):

Thanks :)

OpenStudy (konradzuse):

Wolfram gave me a super weird answer....

OpenStudy (anonymous):

Do not forget to replace u by (1 + sinx) ..